OpenStudy (smurfy14):

A 20 m ladder and a 15 m ladder were leaned against a building. The bottom of the longer ladder was 7 m farther from the building than the bottom of the shorter ladder, but both ladders reached the same distance up the building. Find the distance.

myininaya (myininaya):

i think its 12 let me show you my picture. first let me draw it better k?

OpenStudy (smurfy14):

ok

myininaya (myininaya):

smurfy what do you think?

OpenStudy (smurfy14):

it helps! except the bottom is cutoff..do you think you could repost the bottom part?

myininaya (myininaya):

it just says what h=sqrt{225-81}=sqrt{144}=12

myininaya (myininaya):

i'm painting my nails

OpenStudy (smurfy14):

haha nice and oh ok thanks a bunch!

OpenStudy (anonymous):

if you call the angle between the ladders x then by law of cosines cos(x)=.8 making sin(x)=.6 and therefore the height is 20*.6=12 which i think was already given as an answer yes?

myininaya (myininaya):

np

myininaya (myininaya):

i win satellitle

OpenStudy (anonymous):

i know i am slow

OpenStudy (smurfy14):

satellite73 ya i know but i dont get the whole cos sin thing and this makes more sense

OpenStudy (anonymous):

ahhh no trig for this one?

myininaya (myininaya):

did my without law of cosines so you came up with a differnet way gj

OpenStudy (anonymous):

oh i see your picture. which might in fact be essentially how one proves the law of cosines!

OpenStudy (anonymous):

good work

myininaya (myininaya):

maybe

myininaya (myininaya):

i might check tht out later how to prove that

OpenStudy (anonymous):

kinda quiet tonight. maybe it is break time

OpenStudy (anonymous):

smurfy check the other post you made of this I did it a simpler way

myininaya (myininaya):

if you on break i might catch up with you

OpenStudy (anonymous):

i think that is inevitable.

myininaya (myininaya):

lol

OpenStudy (anonymous):

simpler than the law of cosines version i mean

OpenStudy (anonymous):

oh let me see!

myininaya (myininaya):

i can help 200 people

myininaya (myininaya):

rbsitale did you see my way?

OpenStudy (anonymous):

when you get to 1000 they will send you a bicycle

myininaya (myininaya):

no way

OpenStudy (anonymous):

let me look at yours one sec

OpenStudy (smurfy14):

haha rb ya thats way easier than ur previous way :P

OpenStudy (smurfy14):

*rsvitale not rb lol

OpenStudy (anonymous):

yes myininaya thats the way i did it the second time I wasnt thinking and used law of cosines first

myininaya (myininaya):

oh yay great minds think alike

myininaya (myininaya):

satellitle is brillant too :)

OpenStudy (anonymous):

good I like mathematicians haha

OpenStudy (anonymous):

have you seen any of watchmath's problems he posts some interesting ones here.

myininaya (myininaya):

the calculus dream problem i got close on it i think

OpenStudy (anonymous):

i havent seen that one. I did a few of his others though and they were pretty fun.

myininaya (myininaya):

doing my nails is so hard

myininaya (myininaya):

i want to put an orange pi symbol on all of them after painting them blue

OpenStudy (anonymous):

haha that would be cool

myininaya (myininaya):

don't think it will work

OpenStudy (anonymous):

Ive never seen anyone do that

OpenStudy (anonymous):

take a exacto knife and cut out a small pi stencil then stencil it on

OpenStudy (anonymous):

an*

myininaya (myininaya):

i need more finite space on my nails

myininaya (myininaya):

okay if it works i will post my nails on here later

OpenStudy (smurfy14):

up for helping me with another word problem??

myininaya (myininaya):

i will try to help with one hand

OpenStudy (smurfy14):

ok and ill brb

OpenStudy (smurfy14):

ok im back!

OpenStudy (anonymous):

hope smurfy gets the idea of the solution

OpenStudy (smurfy14):

thanks mark; myin how did you break x^2-48x+560 at the end to get the whole -20 stuff???

myininaya (myininaya):

we had 48=x+560/x i multyplied by x on both sides

OpenStudy (anonymous):

do the law of hypotenuse: c^2 = a^2 + b^2 here the height is C, the base is b , and the opposite sides are 15 and 20 c^2 = 20^2 +(7+x)^2 the other one is c^2 = 15^2 + x^2 now equate them together c^2=c^2 we get 20^2 +(7+x)^2=15^2 + x^2 do the little algebra 20^2 - 15^2 + (7^2 +2*7*x)+x^2=x^2 cancel the x^2 175 + 49 +14x=0 oh here it is 14x=224 x=24/14 x=16 answer

OpenStudy (smurfy14):

in the end you went from x(x-20)-28(x-20)=0 to (x-20)(x-28)=0 how did you know to di that?

OpenStudy (smurfy14):

*do

OpenStudy (smurfy14):

mark its a rectangle so if ur right you need another answer

myininaya (myininaya):

i just factor the long way we have x^2-48x+560=0 if you want you can use the quadratic formula or find factors of 560 that add up to be -48 so we have (x-28)(x-20)=0

OpenStudy (smurfy14):

oh alrighty thanks!

myininaya (myininaya):

lol you like totally get it?

OpenStudy (smurfy14):

ya i do!! :)

myininaya (myininaya):

ok wish me luck on my nails?

OpenStudy (anonymous):

any luck with your answer yet?..lol

OpenStudy (smurfy14):

good luck and have fun! you should attach a picture to this question when youre done and ill check it out :)

myininaya (myininaya):

lol k

OpenStudy (anonymous):

the answer is x= -16.. you can check it by doing the hypotenuse law c^2= a^2 +b^2 c^2= 15^2 +(-16)^2=481 the other one is c^2= 20^2 +(7-16)^2=481 checked